# L-PR: Exploiting LiDAR Fiducial Marker for Unordered Low Overlap Multiview Point Cloud Registration

### v1

@@ Abstract
Point cloud registration is a prerequisite for many applications in computer vision and robotics. Most existing methods focus on pairwise registration of two point clouds with high overlap. Although there have been some methods for low overlap cases, they struggle in degraded scenarios. This paper introduces a novel framework named L-PR, designed to register unordered low overlap multiview point clouds leveraging LiDAR fiducial markers. We refer to them as LiDAR fiducial markers, but they are the same as the popular AprilTag and ArUco markers—thin sheets of paper that do not affect the 3D geometry of the environment. We first propose an improved adaptive threshold marker detection method to provide robust detection results when the viewpoints among point clouds change dramatically. Then, we formulate the unordered multiview point cloud registration problem as a maximum a-posteriori (MAP) problem and develop a framework consisting of two levels of graphs to address it. The first-level graph, constructed as a weighted graph, is designed to efficiently and optimally infer initial values of scan poses from the unordered set. The second-level graph is constructed as a factor graph. By globally optimizing the variables on the graph, including scan poses, marker poses, and marker corner positions, we tackle the MAP problem. We conduct qualitative and quantitative experiments to demonstrate that the proposed method exhibits superiority over competitors in four aspects: (1) registration accuracy, (2) instance reconstruction quality, (3) localization accuracy, and (4) robustness to the degraded scene. To benefit the community, we open-source our method and dataset at https://github.com/yorklyb/LiDAR-SFM.
@@
LiDAR Fiducial Marker; Multiview Point Cloud; Low Overlap; Registration.
@@ I Introduction
Point cloud registration is a fundamental need in computer vision [1, 2] and robotics [3, 4, 5]. Most existing point registration methods [3, 6, 7] focus on pairwise registration of two point clouds with high overlap, which are not robust in practical applications [2]. Some recent learning-based research proposes multiview registration methods [1, 8] and studies low overlap cases [2, 1]. Despite their promising performance on benchmarks, the generalization of learning-based methods to unseen scenarios (i.e., out-of-distribution cases of training data) remains problematic. Moreover, both geometry-based methods [3, 4, 5, 9] and learning-based methods [1, 2, 7] rely on geometric features in the scenes to align point clouds. Hence, these methods struggle in degraded scenes [10], such as those with repetitive structures and weak textures.
@@
In this paper, we develop a novel framework, as shown in Fig. 1(a), to exploit LiDAR fiducial markers for unordered low overlap multiview point cloud registration. As seen, the LiDAR fiducial markers are thin sheets of board/paper attached to other planes without an impact on the environment. Given an unordered set of low overlap point clouds, the proposed method, dubbed L-PR, efficiently registers them in a global frame. To address the unstable marker detection of IFM [20] in the wild, we first develop an adaptive threshold marker detection method that is robust to viewpoint switching. Then, we formulate the multiview point cloud registration problem as a maximum a-posteriori (MAP) problem and develop two-level graphs to tackle it. The first-level graph, constructed as a weighted graph, is designed to efficiently process the unordered set of point clouds and calculate the initial values of poses among point clouds. In particular, the weights represent the pose estimation error of each marker observation, and we obtain the optimal solutions of initial values by finding the shortest path from an anchor scan to each non-anchor scan on the graph. Given the initial values, we design the second-level graph, a factor graph, to resolve the MAP problem by globally optimizing the poses of point clouds, markers, and positions of marker corners. We conduct both qualitative and quantitative experiments to demonstrate the superiority of the proposed method over competitors [1, 3, 22, 5, 23, 4, 15] in four aspects: (1) registration accuracy, (2) instance reconstruction quality, (3) localization accuracy, and (4) robustness to the degraded scene (see Fig. 1(b)).
The contributions of this work are threefold:
  • We design a novel framework, named L-PR, for unordered low overlap multiview point cloud registration leveraging LiDAR fiducial markers.
@@
  • We release the code implementation and datasets at https://github.com/yorklyb/LiDAR-SFM.
@@ II Related Works
LiDAR Fiducial Markers. Given the success of Structure from Motion [13, 15] and visual Simultaneous Localization and Mapping (SLAM) [14, 24] approaches based on visual fiducial markers [11, 12], why do we still need to exploit LiDAR and LiDAR fiducial markers? There are three reasons. (1) Unlike image-based fiducial marker detection, which is sensitive to ambient light, LiDAR fiducial marker detection is robust to unideal illumination conditions [19, 20], such as purely dark or overexposed environments. (2) Unlike image-based fiducial marker detection, which suffers from the rotational ambiguity problem [13, 15, 14], LiDAR fiducial marker pose estimation is not affected by rotational ambiguity [19, 20]. (3) Data captured by LiDARs is irreplaceable by data acquired from cameras due to the unique modality of LiDAR point cloud [25].
@@
Most LiDAR fiducial objects [16, 17, 18] are designed for calibration purposes. Fig. 2(a) shows a typical calibration board, a thick board with holes and/or regions covered by high-intensity materials. A calibration board is commonly placed on a tripod for spatial distinction, making it an additional 3D object that affects the environment. Moreover, while a calibration board can provide fiducials (holes, corners, and high-intensity regions) and it is feasible to assign specific indexes to the fiducials, a calibration board is fundamentally different from fiducial marker systems [11, 12]. The core of fiducial marker systems [11, 12] lies in the elaborately designed encoding-decoding algorithms, leveraging which various complex patterns are generated as depicted in Fig. 2(b). In addition to supporting a larger number of unique IDs, the patterns of fiducial markers are robust against false positive/negative issues. Recently, utilizing the encoding-decoding algorithm of AprilTag 3 [11], the first fiducial marker system for LiDARs, named LiDARTag, is introduced in [19]. Unfortunately, as presented in Fig. 2(c), a LiDARTag is still an additional 3D object impacting the 3D environment. IFM [20] develops a LiDAR fiducial marker system that can be integrated with various visual fiducial marker systems [11, 12]. Furthermore, as seen in Fig. 2(d), the usage/placement of IFM is as convenient as the visual fiducial marker. Nevertheless, the marker detection of IFM is unstable when the viewpoints change dramatically among scans due to the use of a constant threshold to process the intensity image [20]. We develop an improved adaptive threshold marker detection method to address this problem because robust marker detection is the prerequisite for exploiting markers in point cloud registration.
Point Cloud Registration. Geometry-based point cloud registration methods, such as variants of iterative closest point methods [26, 4, 5], RANSAC [9], and Teaser++ [3], mainly focus on designing efficient and robust algorithms for describing geometry and extracting geometric features (e.g., points, lines, and surfaces/normals) to find correspondence between point clouds and align them. Recently, Traj LO [22] proposes to utilize temporal information in addition to geometric information to boost the registration of real-time LiDAR scans. In learning-based point cloud registration methods like Predator [2] and SGHR [1], neural networks are designed to learn features representing the geometry [7, 27, 8, 6], and then learn to utilize these features for registering the point clouds. While these methods demonstrate remarkable performance on benchmarks, they still face challenges in handling degraded scenes, such as repetitive structures and textureless environments. Moreover, most existing methods [26, 4, 5, 7, 6] only apply to high overlap scenarios. Despite the existence of some research [1, 2] on low overlap cases, their generalization to unseen scenes is limited. Inspired by the utilization of visual fiducial markers in camera-based applications [13, 15, 14, 24] to handle degraded scenes and low overlap cases, we propose exploiting LiDAR fiducial markers to enhance multiview point cloud registration.
@@ III-A Adaptive Threshold LiDAR Fiducial Marker Detection
Adaptive Threshold Method. As aforementioned, binarization is applied to the raw intensity image due to the imaging noise (see the zoomed views of Fig. 3). As shown in Fig. 3, the effect is determined by the threshold $\lambda$. In [20], it is found that $\lambda$ can be selected as a constant if the viewpoint does not change drastically. However, with significant changes in the scene, the value of $\lambda$ needs adjustment. This poses a challenge for multiview point cloud registration as the viewpoints of scans can vary dramatically. Therefore, unless an appropriate $\lambda$ is carefully selected for each scan, the LiDAR fiducial marker detection will fail. Take the case shown in Fig. 3 as an example, $\lambda$=12.5 demonstrates the best denoising performance, while other values result in image degradation. Thus, $\lambda$=12.5 is the optimal threshold compared to other values, denoted by $\lambda^{*}$. Unlike [20], which determines $\lambda^{*}$ based on experience or experimentation, we develop an algorithm to automatically seek $\lambda^{*}$.
@@
Given that our focus is on marker detection rather than image denoising, we design Algorithm 1 that utilizes the detection result as the feedback to search for $\lambda^{*}$ automatically. In particular, we design a loop, in which we gradually increase $\lambda$, to determine the threshold value that maximizes the number of detected markers. We experimentally found that it is more efficient to search for $\lambda$ starting from a small value (see Fig. 3). After applying Algorithm 1, the 2D fiducials located on the image binarized with $\lambda^{*}$ are projected into 3D fiducials using the subsequent steps of IFM [20]. Our algorithm resolves the problem of a constant threshold being inapplicable to all scans, especially when LiDAR moves in the wild and scenes undergo significant changes.
@@
Return the image binarized with $\lambda^{*}$.
@@ III-B Problem Formulation
The Singular Value Decomposition (SVD) method [28] is employed to compute $\mathbf{T}^{j,*}_{i}$. We denote the set of measurements, including (1) the corner positions w.r.t. $\{M\}_{j}$, (2) the corner positions w.r.t. $\{F\}_{i}$, and (3) the marker poses w.r.t. $\{F\}_{i}$, as $\mathcal{Z}$.
To register the multiview point clouds, we need to find a globally consistent pose for every scan so that the point clouds can be transformed into a complete point cloud in the global coordinate system $\{G\}$. To achieve this objective, we consider the following variables in this work: (1) the poses of point clouds w.r.t. $\{G\}$, (2) the poses of markers w.r.t. $\{G\}$, and (3) the marker corner positions w.r.t. $\{G\}$. We specify the set of variables as $\Theta$. Finally, a maximum a-posteriori (MAP) inference problem is formulated: given the measurements. $\mathcal{Z}$, the goal is to find the optimal variable set $\Theta^{*}$ that maximizes the posterior probability $P(\Theta\mid\mathcal{Z})$:
@@
In the following, we design a framework consisting of two-level graphs to resolve this MAP problem.
@@ III-C First-level Graph
Eq. (5) indicates the substitution of the result given by SVD back into the right side of Eq. (2). ${e}_{pp}\in\mathbb{R}^{+}$ is employed as the metric to evaluate the quality of pose estimation of the marker in the corresponding point cloud. We define the first scan as the anchor scan. The local coordinate system of the anchor scan is set as the global coordinate system. Namely, $\{F\}_{1}=\{G\}$. We only consider the relative poses between the anchor scan and each non-anchor scan. Although there could be multiple paths from the given scan to the anchor scan (see Fig. 4), we have already saved the information on pose estimation quality by constructing the first-level graph as a weighted graph. Therefore, we adopt Dijkstra’s algorithm [29] to obtain the shortest path. Then, we compute the relative pose along the shortest path iteratively using Eq. (4). Given that the relative pose computation is achieved with the lowest accumulation of $e_{pp}$, the estimation result achieves the highest possible accuracy. Moreover, the search for the optimal path to propagate poses is based merely on the one-dimensional $e_{pp}$, without incorporating any 6-DOF poses or 3D locations.
@@
TABLE I: quantitative comparison with SGHR [1] and Teaser++ [3] on our test dataset

### v2

@@ Abstract
Point cloud registration is a prerequisite for many applications in computer vision and robotics. Most existing methods focus on pairwise registration of two point clouds with high overlap. Although there have been some methods for low overlap cases, they struggle in degraded scenarios. This paper introduces a novel framework dubbed L-PR, designed to register unordered low overlap multiview point clouds leveraging LiDAR fiducial markers. We refer to them as LiDAR fiducial markers, but they are the same as the popular AprilTag and ArUco markers—thin sheets of paper that do not affect the 3D geometry of the environment. We first propose an improved adaptive threshold marker detection method to provide robust detection results when the viewpoints among point clouds change dramatically. Then, we formulate the unordered multiview point cloud registration problem as a maximum a-posteriori (MAP) problem and develop a framework consisting of two levels of graphs to address it. The first-level graph, constructed as a weighted graph, is designed to efficiently and optimally infer initial values of scan poses from the unordered set. The second-level graph is constructed as a factor graph. By globally optimizing the variables on the graph, including scan poses, marker poses, and marker corner positions, we tackle the MAP problem. We conduct both qualitative and quantitative experiments to demonstrate that the proposed method surpasses previous state-of-the-art (SOTA) methods and to showcase that L-PR can serve as a low-cost and efficient tool for 3D asset collection and training data collection. In particular, we collect a new dataset named Livox-3DMatch using L-PR and incorporate it into the training of the SOTA learning-based method, SGHR, which brings evident improvements for SGHR on various benchmarks. We release the open-source code implementation and datasets at https://github.com/yorklyb/L-PR.
@@
LiDAR Fiducial Marker; Multiview Point Cloud; Low Overlap; Registration; Dataset.
@@ I Introduction
Point cloud registration is a fundamental need in computer vision [1, 2] and robotics [3, 4, 5]. Most existing point registration methods [3, 6] focus on pairwise registration of two point clouds with high overlap, which are not robust in practical applications [2]. Some recent learning-based research proposes multiview registration methods [1, 7] and studies low overlap cases [2, 1]. Despite their promising performance on benchmarks, the generalization of learning-based methods to unseen scenarios (i.e., out-of-distribution cases of training data) remains problematic. Moreover, both geometry-based methods [3, 4, 5, 8] and learning-based methods [1, 2] rely on geometric features in the scenes to align point clouds. Hence, these methods struggle in degraded scenes [9], such as those with repetitive structures and weak textures.
Furthermore, training data is crucial for the development of learning-based point cloud registration methods [2, 1], but collecting it is time-consuming and expensive. For instance, in previous dataset collections [6, 21, 22], external sensors such as cameras, IMUs, and GPS are employed to obtain ground truth poses among point clouds. This requires careful multi-sensor calibration and synchronization, which can be time-consuming and labor-intensive.
In this paper, we develop a novel framework, dubbed L-PR (see Fig. 1), to exploit LiDAR fiducial markers for unordered low overlap multiview point cloud registration. L-PR is a new solution for training data collection that only uses a LiDAR sensor. As seen, the LiDAR fiducial markers are thin sheets of board/paper attached to other surfaces without impacting the environment. Specifically, the LiDAR fiducial markers can be printed on paper using standard printers, offering a low-cost, easy-to-use, and reliable alternative compared to previous external sensors.
@@
Given an unordered set of low overlap point clouds, the proposed L-PR efficiently registers them in a global frame. To address the unstable marker detection of IFM [19] in the wild, we first develop an adaptive threshold marker detection method that is robust to viewpoint switching. Then, we formulate the multiview point cloud registration problem as a maximum a-posteriori (MAP) problem and develop two-level graphs to tackle it. The first-level graph, constructed as a weighted graph, is designed to efficiently process the unordered set of point clouds and calculate the initial values of poses among point clouds. In particular, the weights represent the pose estimation error of each marker observation, and we obtain the optimal solutions of initial values by finding the shortest path from an anchor scan to each non-anchor scan on the graph. Given the initial values, we design the second-level graph, a factor graph, to resolve the MAP problem by globally optimizing the poses of point clouds, markers, and positions of marker corners. We conduct both qualitative and quantitative experiments to demonstrate the superiority of the proposed method over competitors [1, 3, 14, 4] and to showcase the diverse applications of L-PR, including 3D asset collection, training data collection, reconstruction of a degraded scene, localization in a GPS-denied environment, and 3D map merging. We release the code implementation and datasets at https://github.com/yorklyb/L-PR.
The contributions of this work are as follows:
  • We design a novel framework, named L-PR, for unordered low overlap multiview point cloud registration leveraging LiDAR fiducial markers. L-PR can serve as an efficient and reliable tool to collect 3D assets and training data for learning-based methods.
  • We collect a new training dataset called Livox-3DMatch using the proposed L-PR. Livox-3DMatch augments the original 3DMatch training data [6] from 14,400 pairs to 17,700 pairs (a 22.91% increase). By training on this augmented dataset, the performance of the SOTA learning-based method SGHR [1] is improved by 2.90% on 3DMatch [6], 4.29% on ETH [21], and 22.72% (translation) / 11.19% (rotation) on ScanNet [22].
@@ II Related Works
LiDAR Fiducial Markers. Given the success of Structure from Motion [12, 14] and visual Simultaneous Localization and Mapping (SLAM) [13, 23] approaches based on visual fiducial markers [10, 11], why do we still need to exploit LiDAR and LiDAR fiducial markers? There are three reasons. (1) Unlike image-based fiducial marker detection, which is sensitive to ambient light, LiDAR fiducial marker detection is robust to unideal illumination conditions [18, 19], such as purely dark or overexposed environments. (2) Unlike image-based fiducial marker detection, which suffers from rotational ambiguity [12, 14, 13] and requires abandoning ambiguous measurements, LiDAR fiducial marker pose estimation is not affected by rotational ambiguity [18, 19] and each marker observation can be utilized. (3) Data captured by LiDARs is irreplaceable by data acquired from cameras due to the unique modality of LiDAR point cloud [24].
Most LiDAR fiducial objects [15, 16, 17] are designed for calibration purposes. Fig. 2(a) shows a typical calibration board, a thick board with holes and/or regions covered by high-intensity materials. A calibration board is commonly placed on a tripod for spatial distinction, making it an additional 3D object that affects the environment. Moreover, while a calibration board can provide fiducials (holes, corners, and high-intensity regions) and it is feasible to assign specific indexes to the fiducials, a calibration board is fundamentally different from fiducial marker systems [10, 11]. The core of fiducial marker systems [10, 11] lies in the elaborately designed encoding-decoding algorithms, leveraging which various complex patterns are generated as depicted in Fig. 2(b). In addition to supporting a larger number of unique IDs, the patterns of fiducial markers are robust against false positive/negative issues. Recently, utilizing the encoding-decoding algorithm of AprilTag 3 [10], the first fiducial marker system for LiDARs, named LiDARTag, is introduced in [18]. Unfortunately, as presented in Fig. 2(c), a LiDARTag is still an additional 3D object impacting the 3D environment. This is undesirable when collecting training data for learning-based point cloud registration methods. IFM [19] develops a LiDAR fiducial marker system that can be integrated with various visual fiducial marker systems [10, 11]. Furthermore, as seen in Fig. 2(d), the usage/placement of IFM is as convenient as the visual fiducial marker. Nevertheless, the marker detection of IFM is unstable when the viewpoints change dramatically among scans due to the use of a constant threshold to process the intensity image [19]. We develop an improved adaptive threshold marker detection method to address this problem because robust marker detection is the prerequisite for exploiting markers in point cloud registration.
@@
Point Cloud Registration. Geometry-based point cloud registration methods, such as variants of iterative closest point methods [4, 5], RANSAC [8], and Teaser++ [3], mainly focus on designing efficient and robust algorithms for describing geometry and extracting geometric features (e.g., points, lines, and surfaces/normals) to find correspondence between point clouds and align them. In learning-based point cloud registration methods like Predator [2] and SGHR [1], neural networks are designed to learn features representing the geometry [25, 26, 27, 7, 6], and then learn to utilize these features for registering the point clouds. While these methods demonstrate remarkable performance on benchmarks [6, 21, 22], they still face challenges in handling degraded scenes, such as repetitive structures and textureless environments. Moreover, most existing methods [28, 4, 5, 27, 6] only apply to high overlap scenarios. Despite the existence of some research [1, 2] on low overlap cases, their generalization to unseen scenes is limited. Inspired by the utilization of visual fiducial markers in camera-based applications [12, 14, 13, 23] to handle degraded scenes and low overlap cases, we propose exploiting LiDAR fiducial markers to enhance multiview point cloud registration.
@@ III-A Adaptive Threshold LiDAR Fiducial Marker Detection
Adaptive Threshold Method. As aforementioned, binarization is applied to the raw intensity image due to the imaging noise (see the zoomed view of Fig. 3). As seen, the effect is determined by the threshold $\lambda$. In [19], it is found that $\lambda$ can be selected as a constant if the viewpoint does not change drastically. However, with significant changes in the scene, the value of $\lambda$ needs adjustment. As an example, consider the case shown in Fig. 3, where three markers are placed. Markers ID 1 and 4 are detectable when $\lambda$=13, while marker ID 3 is detectable when $\lambda$=70. Thus, $\lambda$=13 and $\lambda$=70 are the optimal thresholds compared to other values for markers ID 1 and 4 and marker ID 3, respectively, denoted by $\lambda^{*}$.
@@
Therefore, unless an appropriate $\lambda$ is carefully selected for each scan, and even for each marker, the LiDAR fiducial marker detection will fail. Unlike [19], which determines $\lambda^{*}$ based on experience or experimentation, we develop an algorithm to automatically seek $\lambda^{*}$. Given that our focus is on marker detection rather than image denoising, we design Algorithm 1 that utilizes the detection result as the feedback to search for $\lambda^{*}$ automatically. In particular, the core of the algorithm is to maximize the length of a memory queue for saving detected markers (denoted by $Q$) by gradually increasing $\lambda$. Namely, we aim to detect as many markers as possible by finding the optimal threshold for each marker in the scene. After applying Algorithm 1, the 2D fiducials located on the image binarized with $\lambda^{*}$ are projected into 3D fiducials using the subsequent steps of IFM [19]. Our algorithm addresses the problem of a constant threshold being inapplicable to all scans and markers, especially when LiDAR moves in the wild and scenes undergo significant changes.
@@
Return the image binarized with $\lambda^{*}$ and $Q$.
@@ III-B Problem Formulation
The Singular Value Decomposition (SVD) method [29] is employed to compute $\mathbf{T}^{j,*}_{i}$. We denote the set of measurements, including (1) the corner positions w.r.t. $\{M\}_{j}$, (2) the corner positions w.r.t. $\{F\}_{i}$, and (3) the marker poses w.r.t. $\{F\}_{i}$, as $\mathcal{Z}$. To register the multiview point clouds, we need to find a globally consistent pose for every scan so that the point clouds can be transformed into a complete point cloud in the global coordinate system $\{G\}$. To achieve this objective, we consider the following variables in this work: (1) the poses of point clouds w.r.t. $\{G\}$, (2) the poses of markers w.r.t. $\{G\}$, and (3) the marker corner positions w.r.t. $\{G\}$. We specify the set of variables as $\Theta$. Finally, a maximum a-posteriori (MAP) inference problem is formulated: given the measurements, $\mathcal{Z}$, the goal is to find the optimal variable set $\Theta^{*}$ that maximizes the posterior probability $P(\Theta\mid\mathcal{Z})$:
@@
In the following, we design a framework consisting of two-level graphs to resolve this MAP problem. Inspired by the coarse-to-fine pipeline of SGHR [1], we develop a first-level graph to efficiently and exhaustively determine the relative poses among point clouds, and a second-level graph to globally optimize the variables.
@@ III-C First-level Graph
Eq. (5) indicates the substitution of the result given by SVD back into the right side of Eq. (2). ${e}_{pp}\in\mathbb{R}^{+}$ is employed as the metric to evaluate the quality of pose estimation of the marker in the corresponding point cloud. We define the first scan as the anchor scan. The local coordinate system of the anchor scan is set as the global coordinate system. Namely, $\{F\}_{1}=\{G\}$. We only consider the relative poses between the anchor scan and each non-anchor scan. Although there could be multiple paths from the given scan to the anchor scan (see Fig. 4), we have already saved the information on pose estimation quality by constructing the first-level graph as a weighted graph.
Therefore, we adopt Dijkstra’s algorithm [30] to obtain the shortest path. Then, we compute the relative pose along the shortest path iteratively using Eq. (4). Given that the relative pose computation is achieved with the lowest accumulation of $e_{pp}$, the estimation result achieves the highest possible accuracy. Moreover, the search for the optimal path to propagate poses is based merely on the one-dimensional $e_{pp}$, without incorporating any 6-DOF poses or 3D locations.
@@
III-D Second-level Graph
In this section, we construct a factor graph to optimize the variables, ultimately addressing the MAP inference problem introduced in Eq. (3). Suppose that the $j$-th marker is detected in the $i$-th scan, the following nodes are added to the second graph. The variable nodes include $\mathbf{T}^{j}$ (6-DOF pose of the $j$-th marker w.r.t. $\{G\}$), $\mathbf{T}_{i}$ (6-DOF pose of the $i$-th scan w.r.t. $\{G\}$), and $\{\mathbf{p}^{j,1},\mathbf{p}^{j,2},\mathbf{p}^{j,3},\mathbf{p}^{j,4}\}$ (3D coordinates of the corners of the $j$-th marker w.r.t. $\{G\}$). The factor nodes include $\mathbf{T}^{j}_{i}$ (6-DOF pose of the $j$-th marker w.r.t. $\{F\}_{i}$), $\{^{j}\mathbf{p}^{j,1},\ ^{j}\mathbf{p}^{j,2},\ ^{j}\mathbf{p}^{j,3},\ ^{j}\mathbf{p}^{j,4}\}$ (3D coordinates of the coroners of the $j$-th marker w.r.t. $\{M\}_{j}$), and $\{_{i}\mathbf{p}^{j,1},\ _{i}\mathbf{p}^{j,2},$ ${}_{i}\mathbf{p}^{j,3},$ ${}_{i}\mathbf{p}^{j,4}\}$ (3D coordinates of the corners of the $j$-th marker w.r.t. $\{F\}_{i}$). Note that by adding variables representing the 3D coordinates of the corners, we can also optimize the fiducial localization results at this stage. For clarity, the first stage in Fig. 5 shows the added nodes and edges when a marker is detected in a scan. The graph becomes the one depicted in the second stage of Fig. 5 after integrating all the marker detection results.
Subsequently, we add a prior factor that connects to the first (anchor) point cloud pose node since we set $\{F\}_{1}$ as $\{G\}$. Finally, we add factor nodes representing the relative poses between the anchor scan and each non-anchor scan. Up to this point, the factor graph is completed, as shown in stage 3 of Fig. 5. Following [31, 32], this indicates that the factorization of the posterior probability $P(\Theta\mid\mathcal{Z})$ has been defined:
$P(\Theta\mid\mathcal{Z})=\prod_{k}P^{(k)}(\Theta),$ (6)
We utilize the Levenberg-Marquardt algorithm [32] to solve this problem. The acquisition of initial values is introduced in Section III-C. The noise covariance matrices are determined by the quantitative experiments conducted in [19].
IV Livox-3DMatch Dataset
A common approach [1, 2] to evaluating a learning-based point cloud registration model is to train it on 3DMatch [6] and test it on various benchmarks, including 3DMatch [6], ETH [21], and ScanNet [22]. However, the 3DMatch benchmark is mainly constructed from RGB-D camera captures of indoor scenes [6]. In this work, we collect a new dataset named Livox-3DMatch to enrich the training data for learning-based methods. The enrichment contains two key components. Firstly, the sensor we adopt is a Livox MID-40 LiDAR, which has different sampling patterns compared to the RGB-D camera (see Fig. 6(a) and (b)). Thus, it adds point clouds with new features to the training data. Secondly, we selectively sampled some scenes that are absent or rare in 3DMatch, thereby enriching the scenes in the training data. For example, the valid depth range of an RGB-D camera is usually less than ten meters, making it unsuitable for sampling outdoor scenes. In contrast, we collect some outdoor scenes (see Fig. 6(c)) for Livox-3DMatch, considering that a LiDAR can sample objects a few hundred meters away. Moreover, we collect some challenging cases (see Fig. 6(d)) where the overlapping regions are mainly planes, which are rare in 3DMatch [6]. A more detailed introduction to the scenes in Livox-3DMatch is given in Sections V-B and V-C. In Section V-D, we demonstrate that the proposed Livox-3DMatch can boost the performance of the SOTA learning-based method [1] on various benchmarks [6, 21, 22].
Figure 6: Comparison of 3DMatch and Livox-3DMatch. (a): A random sample from 3DMatch. The point cloud sampled by an RGB-D camera has a regular pattern and less noise. (b): A random sample from Livox-3DMatch. The Livox LiDAR point cloud has an irregular pattern and more noise. (c): An example of an outdoor scene in Livox-3DMatch. (d): A selectively sampled challenging case in which the overlapping regions are mainly planes.
V Experimental Results
In this section, we report the experimental results of seven tests to demonstrate the effectiveness and applications of the proposed method from different perspectives.
V-A Evaluation of the Adaptive Marker Detection Method
Data. We place the Livox MID-40 LiDAR and 69.2 cm $\times$ 69.2 cm ArUco marker(s) in three different scenes, as shown in Fig. 7. The three scenes are: between groups of buildings, in open outdoor areas, and in large indoor parking lots. In each scene, we collect point clouds and test the proposed adaptive threshold method at different relative positions. Results and Analysis. Table I presents the results, where x and y indicate the relative position of the LiDAR in the marker (ID 4) coordinate system. Specifically, in scene Fig. 7(c), there are two markers (ID 4 and ID 1), so the column for $\lambda^{*}$ contains two values. As seen in the table, $\lambda^{*}$ varies significantly as the relative position changes, demonstrating the necessity of the proposed adaptive marker detection method. In addition, the results shown in Fig. 7(c) also illustrate the necessity of the memory queue design, considering that a single $\lambda^{*}$ might not be applicable to all markers in the same scene.
Figure 7: Setup for testing the adaptive threshold marker detection algorithm: (a) between groups of buildings, (b) in open outdoor areas, and (c) in large indoor parking lots.
TABLE I: Demonstration of the necessity of the proposed adaptive marker detection algorithm
Scene x (m) y (m) $\lambda^{*}$ Scene x (m) y (m) $\lambda^{*}$ Scene x (m) y (m) $\lambda^{*}$
Fig. 7(a) 0 4 36 Fig. 7(b) 0 4 7 Fig. 7(c) 4 4 20/6
4 4 9 4 4 8 4 4 8/8
-4 4 6 -4 4 6 -4 4 6/8
0 5 20 0 5 9 0 5 15/10
5 5 8 5 5 8 5 5 8/12
-5 5 7 -5 5 9 -5 5 7/13
0 6 19 0 6 9 0 6 8/13
6 6 10 6 6 6 6 6 11/11
-6 6 11 -6 6 8 -6 6 11/10
0 7 17 0 7 9 0 7 19/20
7 7 8 7 7 8 7 7 15/14
-7 7 9 -7 7 9 -7 7 28/16
0 8 15 0 8 13 0 8 10/16
8 8 9 8 8 11 8 8 11/17
-8 8 10 -8 8 15 -8 8 12/15
0 9 17 0 9 10 0 9 13/11
9 9 8 9 9 12 9 9 20/26
-9 9 9 -9 9 12 -9 9 16/12
0 10 14 0 10 13 0 10 24/28
10 10 11 10 10 10 10 10 10/19
-10 10 12 -10 10 15 -10 10 15/19
V-B Evaluation of Point Cloud Registration Accuracy
Data. Given that the existing point cloud registration benchmarks [6, 21, 22] lack fiducial markers in the scenes, we construct a new dataset, as shown in Fig. 8, with the Livox MID-40. As illustrated in the caption of Fig. 8, our dataset cover various indoor and outdoor scenes. Indoors, multiple 16.4 cm × 16.4 cm AprilTags [10] are positioned in the environment. Outdoors, multiple 69.2 cm × 69.2 cm ArUcos [11] are placed in the environment. Note that since the LiDAR fiducial markers in this work are thin sheets of objects attached to other surfaces, they are almost invisible in the point clouds. This is infeasible if we adopt LiDARTags [18] or calibration boards [15, 16, 17] to provide fiducials. These scenes are challenging due to low overlap, and the overlap rate [2] of each scene is presented in Table II. Competitors and Metrics. Competitors include the state-of-the-art (SOTA) learning-based multiview point cloud registration method, SGHR [1], and the SOTA pairwise geometry-based method, Teaser++ [3]. We employ the root-mean-square errors (RMSEs) [33] as the metric. We also compare the inference time of the methods with an AMD Ryzen 7 5800X CPU. Results and Analysis. The qualitative and quantitative results are presented in Fig. 8 and Table II, respectively. As seen, the proposed method yields the best accuracy and efficiency. SGHR [1] successfully registers the point clouds in scenes 1 to 3 while failing in scenes 4 to 10. Its efficiency is poor on the CPU, while the inference time is around 10s to 20s on an RTX 3090 GPU. Considering that Teaser++ [3] is a pairwise method, we manually select pairs of point clouds with overlap as the inputs (the same scheme adopted in the following tests), yet it still fails in all scenes. The failure cases show SGHR’s limited generalization to unseen scenarios, and Teaser++ struggles with low overlap cases.
TABLE II: quantitative comparison with SGHR [1] and Teaser++ [3] on our dataset